here is my word problem:
Harris Teeter sold 45 tubes of toothpaste, Colgate at $8.50 each and Crest at $9.75 each. In all, $398.75 was taken in for toothpaste. How many of each kind were sold?

Found 2 solutions by nerdybill, angelmidnight:
Answer by nerdybill(7384)

Harris Teeter sold 45 tubes of toothpaste, Colgate at $8.50 each and Crest at $9.75 each. In all, $398.75 was taken in for toothpaste. How many of each kind were sold?
.
Let x = number sold at 8.50
then
45-x = number sold at 9.75
.
8.50x + 9.75(45-x) = 398.75
8.50x + 438.75 - 9.75x = 398.75
8.50x - 9.75x = -40
-1.25x = -40
x = 32 (number sold at $8.50)
.
number sold at $9.75
45-x = 45-32 = 13




Answer by angelmidnight(6)

x+y=45
8.50x+9.75y=398.75
-8.50x+8.50y=-382.5
8.50x+9.75y=398.75
1.25y=16.25
--- ---
1.25 1.25
y=13
x+13=45
x=32
32 Colgate and 13 Crest